# Projection-free computation of robust controllable sets with constrained zonotopes

### v1

@@ 3.2 Inner-approximation of the Pontryagin difference
6:if $D_{ii}\geq 0,\forall i\in\mathbb{N}_{[1:N_{C}]}$ then
7:     $\mathcal{M}^{-}\leftarrow(G_{C}D,c_{C}-c_{S},A_{C}D,b_{C})$
8:else
9:     $\mathcal{M}^{-}\leftarrow\emptyset$
10:end if
Algorithm 1 summarizes the procedure described in Theorems 1 and 2 to inner-approximate $\mathcal{C}\ominus\mathcal{S}$ when the minuend $\mathcal{C}$ is full-dimensional (Assumption 1). For subtrahends that are ellipsoids, zonotopes, or convex unions of symmetric intervals, Step 5 in Algorithm 1 is available in closed form (see Corollary 1), and Algorithm 1 is optimization-free, i.e., it does not require convex optimization solvers.
3.3 Relation of Algorithm 1 with existing two-stage approach for zonotopic subtrahend
For a zonotopic subtrahend $\mathcal{S}$, we show that the optimization-free Algorithm 1 is closely related to the optimization-based two-stage approach [9] (see Section 2.2). From (1d), $\mathcal{S}=G_{S}\mathcal{S}_{0}+c_{S}$ where $\mathcal{S}_{0}$ is the unit $\ell_{\infty}$-norm ball.
Consider the matrix least norm problem [20, Ex. 16.2], which is similar to (12) where the objective is now the Frobenius norm $\|\Gamma\|_{F}\triangleq\sqrt{\sum_{i=1}^{N_{C}}\sum_{j=1}^{N_{S}}\Gamma_{ij}^{2}}$,
@@ 3.3 Relation of Algorithm 1 with existing two-stage approach for zonotopic subtrahend
When $\mathcal{C}$ is full-dimensional, the optimal solution of (41) is available in closed-form, $\Gamma^{\ast}=[G_{C};A_{C}]^{\dagger}[G_{S};0_{M_{C}\times N_{S}}]$, by Proposition 1 and [20, Ex. 16.2]. Observe that $\Gamma^{\ast}=\Gamma G_{S}$ for $\Gamma$ prescribed by Theorem 2. Then, we can recover $D$ prescribed by (37) in Corollary 1 using (26) in Theorem 1, where $\rho_{\mathcal{S}_{0}}(\nu)=\|\nu\|_{1}$ instead of (5c).
The condition $D_{ii}\geq 0$ is required both in the two-stage approach [9] (see (12)) and in Step 6 of Algorithm 1. Thus, the expression for $D_{ii}$ obtained by Algorithm 1 is identical to that in (13), and Algorithm 1 and [9] differ only in the choice of $\Gamma$ (or specifically, the choice of objective in (12)) for a zonotopic subtrahend. In other words, Algorithm 1 may be viewed as a generalization of [9] to handle any symmetric, convex, and compact subtrahend. Moreover, thanks to Theorem 2, Algorithm 1 has some computational advantage over [9], as discussed further in Section 6.3.
@@ 4 Outer-approximation of Pontryagin difference and sufficient conditions for exactness
In this section, we first address Problem 2, and then use its solution to address the outer-approximation part of Problem 1. We also provide sufficient conditions under which all proposed approximations are exact. We conclude this section with a discussion of implementation considerations for the proposed algorithms.
4.1 Constructing an outer-approximating convex polyhedron for a given constrained zonotope
@@
2:Convex polyhedron $\mathcal{P}=\{x:Hx\leq k\}\supseteq\mathcal{C}$
@@
5:$k\leftarrow 1_{(2N_{C})\times 1}-[V;-V][-c_{C};b_{C}]$
We address Problem 2 using Algorithm 2 that computes a convex, polyhedral outer-approximation of a constrained zonotope.
@@
Given a full-dimensional constrained zonotope $\mathcal{C}$ (Assumption 1), Algorithm 2 computes an convex, polyhedral, outer-approximation of $\mathcal{C}$ in the form of (1a) with at most $2N_{C}$ linear constraints.
See Appendix A.2 for the proof of Proposition 2. The key insight used in Proposition 2 is that we can express $\mathcal{C}$ in (1b) as the $1$-sublevel set of the optimal value function of a linear program that is always feasible for a full-dimensional $\mathcal{C}$ (see (83) in Appendix A.2). Then, we use strong duality to obtain an outer-approximating convex polyhedron $\mathcal{P}$, characterized by $2N_{C}$ feasible solutions (42) to the corresponding dual problem.
While [6, Prop. 3] provides an exact convex polytopic representation of $\mathcal{C}$ using lifted zonotopes, it may require a combinatorial number of hyperplanes, and may not be practical for large $n$ and/or $N_{C}$. Instead, Algorithm 2 outer-approximates $\mathcal{C}$ with a convex polyhedron $\mathcal{P}$ that has at most $2N_{C}$ halfspaces, and can be efficiently computed for $\mathcal{C}$ with large $n$ and $N_{C}$ without relying on convex optimization solvers. Redundant inequalities in $\mathcal{P}$ may be removed via linear programming [4], if desired.
Additional assumptions on $\mathcal{C}$ may be required to ensure that $\mathcal{P}$ is bounded (and, thereby a convex polytope), and representable as a constrained zonotope using [6, Thm. 1]. We provide one such assumption in Section 4.3.
@@ 4.2 Outer-approximation of the Pontryagin difference
Next, we address the outer-approximation part of Problem 1. We use Algorithm 2, (8), and (7d) to compute $\mathcal{M}^{+}$, an outer-approximation of the Pontryagin difference $\mathcal{C}\ominus\mathcal{S}$, and summarize the approach in Algorithm 3.
@@
3:Compute a convex polyhedron $\mathcal{P}\supseteq\mathcal{C}$ using Algorithm 2
4:Compute a convex polyhedron $\mathcal{M}^{+}_{\text{p}}=\mathcal{P}\ominus\mathcal{S}$ via (8)
5:Compute $\mathcal{M}^{+}\leftarrow(\mathcal{C}-c_{S})\cap\mathcal{M}^{+}_{\text{p}}$ using (7d)
@@
For any full-dimensional $\mathcal{C}$ and any convex and compact subtrahend $\mathcal{S}$ that is symmetric about any $c_{S}\in\mathbb{R}^{n}$, Algorithm 3 returns $\mathcal{M}^{+}\supseteq\mathcal{C}\ominus\mathcal{S}$.
@@
By Proposition 2 and (8), $\mathcal{M}^{+}_{\text{p}}$ constructed in Step 4 of Algorithm 3 outer-approximates $\mathcal{C}\ominus\mathcal{S}$. Since $\mathcal{S}$ is symmetric about $c_{S}$, $(\mathcal{C}-c_{S})$ is also an outer-approximation of $\mathcal{C}\ominus\mathcal{S}$ by (6d). We obtain the outer-approximating constrained zonotope $\mathcal{M}^{+}$ by intersecting these outer-approximations in Step 5 using (7d). ∎
All steps of Algorithm 3 are available in closed-form, when the support function of $\mathcal{S}$ is known in closed-form. Similarly to Algorithm 1, Algorithm 3 is optimization-free, i.e., it does not require convex optimization solvers, for subtrahends that are ellipsoids, zonotopes, or convex unions of symmetric intervals (see (2.1)).
@@ 4.3 Sufficient conditions for exactness
Now, we state a sufficient condition under which the approximations of the Pontryagin difference computed by Algorithms 1 and 3, and the convex polyhedral outer-approximation of a constrained zonotope computed by Algorithm 2 are exact.
Assumption 2.
The minuend $\mathcal{C}$ has a MinRow representation (1b) such that $n+M_{C}=N_{C}$.
When Assumption 2 holds, $[G_{C};A_{C}]$ is square and invertible by Proposition 1, since Assumption 2 implies Assumption 1.
@@
(Sufficient conditions for exactness of polyhedral cover) Under Assumption 2, Algorithm 2 computes a convex polytope $\mathcal{P}=\mathcal{C}$.
@@
(Sufficient conditions for exactness of Pontryagin difference) Under Assumption 2, the approximations $\mathcal{M}^{-}$ and $\mathcal{M}^{+}$ computed using Algorithms 1 and 3 satisfy $\mathcal{M}^{-}=\mathcal{M}=\mathcal{C}\ominus\mathcal{S}=\mathcal{M}^{+}$.
See Appendices A.3 and A.4 for the proofs. Under Assumption 2, $\mathcal{M}^{+}_{\text{p}}$ computed in Step 4 of Algorithm 3 is $\mathcal{C}\ominus\mathcal{S}$ by Proposition 4. Consequently, we can skip Step 5 of Algorithm 3, and directly compute a constrained zonotope representation of $\mathcal{M}$ from the convex polytope $\mathcal{M}^{+}_{\text{p}}$.
Algorithm 13 together address Problem 1, as shown by Theorems 12, and Propositions 35. Algorithm 2 addresses Problem 2, as shown by Propositions 2 and 4.
@@ 4.4 Implementation considerations
The use of pseudoinverse $[G_{C};A_{C}]^{\dagger}$ in Algorithms 12, and 3 was motivated by providing closed-form expressions for $D_{ii}$ (26) and $v_{i}$ (42). However, the computation of $[G_{C};A_{C}]^{\dagger}$ can be computationally expensive for large $N_{C}$ and $M_{C}$. In practice, it suffices to compute a minimum norm solution of systems of linear equations — a solution $\Gamma$ in (25) for Step 4 of Algorithm 1, a solution $\xi$ to $A_{C}\xi=b_{C}$ in Step 6 of Algorithm 1, and a solution $V^{\top}$ to $[G_{C};A_{C}]^{\top}V^{\top}=I_{N_{C}}$ for Step 3 of Algorithm 2 (where $V[G_{C};A_{C}]$ is later normalized row-wise in $\ell_{1}$-norm).
We can also compute a minimum norm solution without computing the psuedoinverse explicitly via QR factorization or complete orthogonal decomposition [20, Ch. 12]. Existing algorithms can also exploit sparsity [20, Ch. 12.3]. In Sections 6 and 7, our MATLAB implementation of Algorithms 12, and 3 utilizes lsqminnorm to compute minimum norm solutions and uses sparse matrices for computational efficiency.
@@ 5 Inner-approximation of robust controllable sets
We now address Problem 3 by inner-approximating the $T$-step RC set using Algorithm 1 and the set recursion (2.3). We consider both cases described in Problem 3, characterize the representation complexity of the computed inner-approximations, and show that their representation complexities grow linearly with $T$.
Throughout this section, we will assume that the input set $\mathcal{U}_{t}$ and the goal set $\mathcal{G}$ are convex polytopes, hence, representable as constrained zonotopes, and the additive disturbance set $\mathcal{W}_{t}$ is a convex and compact set that is symmetric about any $c_{W}\in\mathbb{R}^{p}$.
@@ 5.1 Convex polyhedral $\mathcal{X}_{t}$ and invertible $A_{t}$
Given a finite horizon $T\in\mathbb{N}$, we consider the case where $A_{t}$ matrices of the linear system (16) are invertible, and state constraints $\mathcal{X}_{t}$ are polyhedra for all $t\in\mathbb{N}_{[0:T-1]}$. Since $\mathcal{X}_{t}$ can be unbounded, they may not be representable as constrained zonotopes. However, we can still inner-approximate the RC sets as constrained zonotopes using (7d), as shown in this section.
For all $t\in\mathbb{N}_{[0:T-1]}$, we break down the set recursion (2.3) to compute the $T$-step RC set into four steps:
@@
The recursion (5.1) is initialized with a constrained zonotope $\mathcal{K}_{T}=\mathcal{G}$. We use Theorem 2 to compute a constrained zonotope $\mathcal{K}_{t,\text{inner}}^{\text{interim,1}}$ in (43a). Then, we compute (43b)–(43d) exactly using (2.2). Thus, for all $t\in\mathbb{N}_{[0:T-1]}$, the sets $\mathcal{K}_{t},\mathcal{K}_{t,\text{inner}}^{\text{interim,1}},\mathcal{K}_{t}^{\text{interim,2}}$, and $\mathcal{K}_{t}^{\text{interim,3}}$ are constrained zonotopes, and $\mathcal{K}^{-}=\mathcal{K}_{0}$ is an inner-approximation of the $T$-step RC set.
Set $\mathcal{S}$ Eq. no. $N_{S}$ $M_{S}$
$\mathcal{K}_{t,\text{inner}}^{\text{interim,1}}$ (43a) $N_{K,t+1}$ $M_{K,t+1}$
$\mathcal{K}_{t}^{\text{interim,2}}$ (43b) $N_{K,t+1}+N_{U}$ $M_{K,t+1}+M_{U}$
$\mathcal{K}_{t}^{\text{interim,3}}$ (43c) $N_{K,t+1}+N_{U}$ $M_{K,t+1}+M_{U}$
$\mathcal{K}_{t}$ (43d) $N_{K,t+1}+N_{U}+N_{X}$ $M_{K,t+1}+M_{U}+N_{X}$
Table 2: Representation complexity (see Definition 1) for various sets involved in computing an inner-approximation to the $T$-step RC set using (5.1), where $\mathscr{C}(\mathcal{K}_{t+1})=(n,N_{K_{t+1}},M_{K_{t+1}})$ and $\mathscr{C}(\mathcal{U})=(m,N_{U},M_{U})$, and $\mathcal{X}$ is characterized by $N_{X}$ hyperplanes. Observe that the representation complexity grows by $(0,N_{U}+N_{X},M_{U}+N_{X})$ with each step of the recursion.
Table 2 describes the representation complexity of various constrained zonotopes involved at each step of (5.1). For ease of discussion, we assume that $\mathcal{X}_{t},\mathcal{U}_{t}$, and $\mathcal{W}_{t}$ are time-invariant, i.e., $\mathcal{X}_{t}=\mathcal{X}$, $\mathcal{U}_{t}=\mathcal{U}$, and $\mathcal{W}_{t}=\mathcal{W}$, for all $t\in\mathbb{N}_{[0:T]}$. Additionally, we assume that $\mathcal{X}$ is characterized by $N_{X}$ hyperplanes. Given representation complexities $\mathscr{C}(\mathcal{K}_{t+1})=(n,N_{K_{t+1}},M_{K_{t+1}})$ and $\mathscr{C}(\mathcal{U})=(m,N_{U},M_{U})$, the rows of Table 2 follow from Theorem 2, (7b), (7a), and (7d), respectively.
With $\mathscr{C}(\mathcal{G})=(n,N_{G},M_{G})$, the representation complexity of the inner-approximation of the $T$-step RC set is
$\mathscr{C}(\mathcal{K}^{-})=(n,\ N_{G}+T(N_{U}+N_{X}),\ M_{G}+T(M_{U}+N_{X})).$ (44)
Observe that the representation complexity of $\mathcal{K}^{-}$ does not depend on the disturbance set $\mathcal{W}$, due to Theorem 2, and grows linearly with $T$.
5.2 Convex polytopic $\mathcal{X}_{t}$
We now consider the case where the state constraints $\mathcal{X}_{t}$ are convex polytopes for all $t\in\mathbb{N}$, and are representable using constrained zonotopes. Unlike Section 5.1, we no longer assume that $A_{t}$ is invertible.
Similarly to (5.1), we separate the set recursion (2.3) into three steps performed for all $t\in\mathbb{N}_{[0:T-1]}$:

### v2

@@ 3.2 Inner-approximation of the Pontryagin difference
6:$\mathcal{M}^{-}\leftarrow\begin{cases}\begin{array}[]{ll}(G_{C}D,c_{C}-c_{S},A_{C}D,b_{C}),&\quad\min\limits_{i\in\mathbb{N}_{[1:{N_{C}}]}}D_{ii}\geq 0,\\ \emptyset,&\quad\text{otherwise}\end{array}.\end{cases}$
Algo. 3 summarizes the procedure described in Prop. 2 to inner-approximate $\mathcal{C}\ominus\mathcal{S}$ when the minuend $\mathcal{C}$ is full-dimensional. For subtrahends that are ellipsoids, zonotopes, or convex unions of symmetric intervals, Step 5 in Algo. 3 is available in closed form (see Corr. 1), and Algo. 3 is optimization-free, i.e., it does not require convex optimization solvers.
Remark 2.
We can also use a “template” constrained zonotope $\mathcal{C}^{-}\subseteq\mathcal{C}$ when available and compute $\mathcal{C}^{-}\ominus\mathcal{S}\subseteq\mathcal{C}\ominus\mathcal{S}$ using Thm. 1. Such an approach may help in overcoming conservativeness in certain cases, e.g., $\mathcal{M}^{-}$ described in Thm. 1 is a zonotope for zonotopes $\mathcal{C},\mathcal{S}$, even when $\mathcal{C}\ominus\mathcal{S}$ is known to be a constrained zonotope [9, 6].
3.3 Relation of Algo. 3 with existing two-stage approach for zonotopic subtrahend
For a zonotopic subtrahend $\mathcal{S}$, the optimization-free Algo. 3 is closely related to the optimization-based two-stage approach [9] (see Sec. 2.2).
Let $\mathcal{S}=G_{S}\mathcal{S}_{0}+c_{S}$ where $\mathcal{S}_{0}$ is the unit $\ell_{\infty}$-norm ball (3a). Consider the matrix least norm problem [21, Ex. 16.2], which is similar to (12) where the objective is now the Frobenius norm $\|\Gamma\|_{F}\triangleq\sqrt{\sum_{i=1}^{{N_{C}}}\sum_{j=1}^{{N_{S}}}\Gamma_{ij}^{2}}$,
@@
When $\mathcal{C}$ is full-dimensional, the optimal solution of (47) is available in closed-form, $\Gamma^{\ast}=[G_{C};A_{C}]^{\dagger}[G_{S};0_{{M_{C}}\times{N_{S}}}]$, by Prop. 1.2 and [21, Ex. 16.2]. Observe that $\Gamma^{\ast}=\Gamma G_{S}$ for $\Gamma$ prescribed by Prop. 2. Then, we can recover $D$ prescribed by (42) in Corr. 1 using (27) in Thm. 1, where $\rho_{\mathcal{S}_{0}}(\nu)=\|\nu\|_{1}$ instead of (5c).
The condition $D_{ii}\geq 0$ for every $i\in\mathbb{N}_{[1:{N_{C}}]}$ is required in the two-stage approach [9] (see (12)) and in Step 6 of Algo. 3. Thus, Algo. 3 and [9] differ only in the choice of $\Gamma$ (or specifically, the choice of objective in (12)) for a zonotopic subtrahend. In other words, Algo. 3 may be viewed as a generalization of [9] for symmetric, convex, and compact subtrahends, and does not require an optimization solver for certain subtrahends (Corr. 1). We compare these approaches in various examples in Sec. 6.
@@ 4 Outer-approximation of Pontryagin difference and sufficient conditions for exactness
In this section, we first address Prob. 2, and then use its solution to address the outer-approximation part of Prob. 1. We also provide sufficient conditions under which all proposed approximations are exact. We conclude this section with a discussion of implementation considerations for the proposed algorithms.
4.1 Outer-approximating convex polyhedron for a given constrained zonotope
@@
2:Convex polyhedron $\mathcal{P}=\{x\mid Hx\leq k\}\supseteq\mathcal{C}$
@@
5:$k\leftarrow 1_{2{N_{C}}\times 1}-[V;-V][-c_{C};b_{C}]$
We now address Prob. 2 using Algo. 4.
@@
Given a full-dimensional constrained zonotope $\mathcal{C}$, Algo. 4 computes a convex, polyhedral, outer-approximation of $\mathcal{C}$ with at most $2{N_{C}}$ linear constraints.
See Sec. A.2 for the proof. We used the observation that a full-dimensional $\mathcal{C}$ may be expressed as the $1$-sublevel set of the optimal value function of a feasible linear program. Using strong duality, we obtain an outer-approximating convex polyhedron $\mathcal{P}$, characterized by $2{N_{C}}$ feasible solutions (48) to the corresponding dual problem.
While [5, Prop. 3] provides an exact H-Rep polytope representation of $\mathcal{C}$ using lifted zonotopes, it may require a combinatorial number of hyperplanes, and may not be practical for large $n$ and/or ${N_{C}}$. Instead, Algo. 4 outer-approximates $\mathcal{C}$ with a convex polyhedron $\mathcal{P}$ that has at most $2{N_{C}}$ halfspaces, and can be efficiently computed for $\mathcal{C}$ with large $n$ and ${N_{C}}$ without relying on convex optimization solvers. Redundant inequalities in $\mathcal{P}$ may be removed via linear programming [24], if desired.
Remark 3.
(H-Rep outer-approximation) We can modify Algo. 4 in the following ways:
1) Always return a H-Rep polytope $\mathcal{P}\cap\{x\mid l\leq x\leq u\}$ with $l,u$ computed in (24), where ${L_{P}}=2({N_{C}}+n)$.
2) Reduce approximation error by intersecting $\mathcal{P}$ computed by Algo. 4 with supporting hyperplanes (4) evaluated along template directions (ray shooting) [10].
Both modifications incur additional computational cost.
@@ 4.2 Outer-approximation of the Pontryagin difference
We now address the outer-approximation part of Prob. 1.
@@
3:Compute a convex polyhedron $\mathcal{P}\supseteq\mathcal{C}$ using Algo. 4.
4:Compute a convex polyhedron $\mathcal{M}^{+}_{\text{p}}=\mathcal{P}\ominus\mathcal{S}$ via (8).
5:Compute $\mathcal{M}^{+}\leftarrow(\mathcal{C}-c_{S})\cap\mathcal{M}^{+}_{\text{p}}$ using (7d) .
@@
For a full-dimensional, constrained zonotopic minuend $\mathcal{C}$ and a convex and compact subtrahend $\mathcal{S}\subset\mathbb{R}^{n}$ that is symmetric about any $c_{S}\in\mathbb{R}^{n}$, Algo. 5 returns a constrained zonotope $\mathcal{M}^{+}\supseteq\mathcal{C}\ominus\mathcal{S}$.
@@
By Prop. 3 and (8), $\mathcal{M}^{+}_{\text{p}}$ constructed in Step 4 of Algo. 5 outer-approximates $\mathcal{C}\ominus\mathcal{S}$. Since $\mathcal{S}$ is symmetric about $c_{S}$, $(\mathcal{C}-c_{S})$ is also an outer-approximation of $\mathcal{C}\ominus\mathcal{S}$ by (6d). We obtain the outer-approximating constrained zonotope $\mathcal{M}^{+}$ by intersecting these outer-approximations in Step 5 using (7d). ∎
All steps of Algo. 5 are available in closed-form, when the support function of $\mathcal{S}$ is known in closed-form. Similarly to Algo. 3, Algo. 5 is optimization-free, i.e., it does not require convex optimization solvers, when $\mathcal{S}$ is an ellipsoid, a zonotope, or a convex unions of symmetric intervals (see (5)).
@@ 4.3 Sufficient conditions for exactness
We show that the proposed approximations become exact when using an Invertible representation.
@@
(Sufficient conditions for exactness of polyhedral cover) For an Invertible representation $\mathcal{C}$, Algo. 4 computes a H-Rep polytope $\mathcal{P}=\mathcal{C}$ that is identical to the H-Rep polytope in Prop 1.5.
@@
(Sufficient conditions for exactness of Pontryagin difference) For an Invertible representation $\mathcal{C}$, Algo. 3 and 5 provide approximations $\mathcal{M}^{-}$ and $\mathcal{M}^{+}$ that satisfy $\mathcal{M}^{-}=\mathcal{M}=\mathcal{C}\ominus\mathcal{S}=\mathcal{M}^{+}$.
See Sec. A.3 and A.4 for the proofs. The exactness results (Prop. 5 and 6) may be attributed to Prop. 1.5. Also, for $\mathcal{C}$ in an Invertible representation, Step 4 of Algo. 5 computes a H-Rep polytope $\mathcal{M}^{+}_{\text{p}}=\mathcal{C}\ominus\mathcal{S}$ by Prop. 5. Consequently, instead of Step 5, we can use Algo. 2 to compute $\mathcal{M}^{+}$ directly from $\mathcal{M}_{p}^{+}$.
Algo. 35 address Prob. 1 by Thm. 1 and Prop. 246. Algo. 4 addresses Prob. 2 by Prop. 35.
@@ 4.4 Implementation considerations
The use of pseudoinverse $[G_{C};A_{C}]^{\dagger}$ in Algo. 34, and 5 was motivated by providing closed-form expressions for $D_{ii}$ (27) and $v_{i}$ (48). However, the computation of $[G_{C};A_{C}]^{\dagger}$ can be computationally expensive for large ${N_{C}}$ and ${M_{C}}$. In practice, it suffices to compute a minimum norm solution of systems of linear equations — a solution $\Gamma$ in (26) for Step 4 of Algo. 3, a solution $\xi$ to $A_{C}\xi=b_{C}$ in Step 6 of Algo. 3, and a solution $V^{\top}$ to $[G_{C};A_{C}]^{\top}V^{\top}=I_{{N_{C}}}$ for Step 3 of Algo. 4 (where $V[G_{C};A_{C}]$ is later normalized row-wise in $\ell_{1}$-norm).
We can use QR factorization or complete orthogonal decomposition to compute a minimum norm solution without explicitly computing the psuedoinverse [21, Ch. 12]. Existing algorithms can also exploit sparsity [21, Ch. 12.3]. In Sec. 6 and 7, our MATLAB implementation of Algo. 34, and 5 utilizes lsqminnorm to compute minimum norm solutions and uses sparse matrices for computational efficiency.
@@ 5 Inner-approximation of robust controllable sets
We now address Prob. 3 by inner-approximating the $T$-step RC set using Algo. 3 and the set recursion (21). We consider both cases described in Prob. 3, characterize the representation complexity of the computed inner-approximations, and show that their representation complexities grow linearly with $T$.
Throughout this section, we will assume that the input set $\mathcal{U}_{t}$ and the goal set $\mathcal{G}$ are polytopes, hence, representable as constrained zonotopes, and the additive disturbance set $\mathcal{W}_{t}$ is a convex and compact set that is symmetric about any $c_{W}\in\mathbb{R}^{p}$. Also, we assume that the sets $\mathcal{K}_{t}$ computed in (21) are full-dimensional for every $t$.
@@ 5.1 Convex polyhedral $\mathcal{X}_{t}$ and invertible $A_{t}$
Set Eq. no. ${M}$ $\mathscr{O}$
$\mathcal{K}_{t,\text{inner}}^{\text{interim,1}}$ (49a) ${M_{\mathcal{K}_{t+1}}}$ $\mathscr{O}_{{\mathcal{K}_{t+1}}}$
$\mathcal{K}_{t}^{\text{interim,2}}$ (49b) ${M_{\mathcal{K}_{t+1}}}+{M_{B\mathcal{U}}}$ $\mathscr{O}_{{\mathcal{K}_{t+1}}}+\mathscr{O}_{{B\mathcal{U}}}$
$\mathcal{K}_{t}^{\text{interim,3}}$ (49c)
$\mathcal{K}_{t}$ (49d) ${M_{\mathcal{K}_{t+1}}}+{M_{B\mathcal{U}}}+{L_{X}}$
Table 2: Representation complexity (see Defn. 1) for various sets involved in computing an inner-approximation to the $T$-step RC set using (49), where $\mathscr{C}(\mathcal{K}_{t+1})=({M_{\mathcal{K}_{t+1}}},\mathscr{O}_{{\mathcal{K}_{t+1}}})$ and $\mathscr{C}(B\mathcal{U})=({M_{B\mathcal{U}}},\mathscr{O}_{{B\mathcal{U}}})$, and $\mathcal{X}$ is characterized by ${L_{X}}$ hyperplanes. Observe that the representation complexity grows by $({M_{BU}}+{L_{X}},\mathscr{O}_{{B\mathcal{U}}})$ with each step of the recursion.
Given a finite horizon $T\in\mathbb{N}$, we consider the case where $A_{t}$ in (16) are invertible, and state constraints $\mathcal{X}_{t}$ are polyhedra for all $t\in\mathbb{N}_{[0:T-1]}$. Since $\mathcal{X}_{t}$ can be unbounded, they may not be representable as constrained zonotopes. However, we can still inner-approximate the RC sets as constrained zonotopes using (7d).
For all $t\in\mathbb{N}_{[0:T-1]}$, we break down the set recursion (21) to compute the $T$-step RC set into four steps:
@@
The recursion (49) is initialized with a constrained zonotope $\mathcal{K}_{T}=\mathcal{G}$. We use Prop. 2 to compute a constrained zonotope $\mathcal{K}_{t,\text{inner}}^{\text{interim,1}}$ in (49a). Then, we compute (49b)–(49d) exactly using (7). Thus, for all $t\in\mathbb{N}_{[0:T-1]}$, the sets $\mathcal{K}_{t},\mathcal{K}_{t,\text{inner}}^{\text{interim,1}},\mathcal{K}_{t}^{\text{interim,2}}$, and $\mathcal{K}_{t}^{\text{interim,3}}$ are constrained zonotopes, and $\mathcal{K}^{-}=\mathcal{K}_{0}$ is an inner-approximation of the $T$-step RC set.
Table 2 describes the representation complexity of various constrained zonotopes involved at each step of (49). For ease of discussion, we assume that $\mathcal{X}_{t},\mathcal{U}_{t}$, and $\mathcal{W}_{t}$ are time-invariant, i.e., $\mathcal{X}_{t}=\mathcal{X}$, $\mathcal{U}_{t}=\mathcal{U}$, and $\mathcal{W}_{t}=\mathcal{W}$, for all $t\in\mathbb{N}_{[0:T]}$. Additionally, we assume that $\mathcal{X}$ is characterized by ${L_{X}}$ hyperplanes. Given representation complexities $\mathscr{C}(\mathcal{K}_{t+1})=({M_{\mathcal{K}_{t+1}}},\mathscr{O}_{{\mathcal{K}_{t+1}}})$ and $\mathscr{C}(B\mathcal{U})=({M_{B\mathcal{U}}},\mathscr{O}_{{B\mathcal{U}}})$, the rows of Table 2 follow from Prop. 2, (7b), (7a), and (7d), respectively. With $\mathscr{C}(\mathcal{G})=({M_{\mathcal{G}}},\mathscr{O}_{{\mathcal{G}}})$, the representation complexity of the inner-approximation of the $T$-step RC set is
$\mathscr{C}(\mathcal{K}^{-})=({M_{\mathcal{G}}}+T({M_{B\mathcal{U}}}+{L_{X}}),\mathscr{O}_{{\mathcal{G}}}+T\mathscr{O}_{{B\mathcal{U}}}).$ (50)
Observe that the representation complexity of $\mathcal{K}^{-}$ does not depend on the disturbance set $\mathcal{W}$ due to Prop. 2, and grows linearly with $T$.
5.2 Polytopic $\mathcal{X}_{t}$
We now consider the case where the state constraints $\mathcal{X}_{t}$ are polytopes for all $t\in\mathbb{N}$, and admit a constrained zonotope representation with $\mathscr{C}(\mathcal{X}_{t})=(L_{\mathcal{X}_{t}},1)$ when using Algo. 2. Unlike Sec. 5.1, we no longer assume that $A_{t}$ is invertible.
Similarly to (49), we separate the set recursion (21) into three steps performed for all $t\in\mathbb{N}_{[0:T-1]}$: